McDonalds big mac contains 44g of carbohydrates, 36g of fat and 37 g of protein. what is the percentage of energy derived from each one (carbohydrates, fat, and protein)?

please show all work.
Answer #1

A healthy man in a day will need 2500kcal /day as a standard recommended intake

Given

Carbohydrates :44 g

(As 1 g of carbohydrate =4calories, then 44g=176 calories _44×4 )

Divide it by total calories and multiply by 100

=176÷2500×100

=7.04% or 7 % of carbohydrates

Protein:37g

(As1 g of protein =4calories, then 37g=148 calorie )

Divide it by total calories and multiply by 100

=148÷2500×100

=5.92% or 6% of protein

Fats:39 g

(As 1g fat =9 g ,then 39g=351 calories )

Divide it by total kcal and multiply by 100

=351÷2500×100

=14.04% or 14% of fat